[Brahms-l] Shift report 20040116 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Fri Jan 16 2004 - 15:59:57 EST
------------------------------------------------------------
Date: Fri Jan 16 15:59:57 2004, Shift: 08:00-16:00
Supervisor: OR, with: 
Log book#15 pages 34 - 34

Shift Summary:
------------------------------------------------------------
1. Machine development for entire shift.

2. Pedestals: run 9710 - non-TPC pedestal run, run 9711 - TPC pedestal run.
   Data: run 9712 FS@12 with D1-861B, MRS@90 with D5=2255A - 129k tr. 6, 40k tr. 3.

3. Some problems with pii3, but Hiro fixed. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 16 16:00:34 2004

This archive was generated by hypermail 2.1.8 : Fri Jan 16 2004 - 16:00:54 EST